A population consists of value (2,3,5). Consider all possible samples of size n=2 that can be drawn with resplacement from this population

1
Expert's answer
2022-05-24T23:02:35-0400

The number of possible samples which can be selected with replacement is

"N^n=3^2=9."

All the possible samples of sizes n=2 wich can be drawn with replacement from the population:

"\{(2,2),(2,3),(2,5),(3,2),(3,3),
\\(3,5),(5,2),(5,3),(5,5)\}."